OpenStudy (kkbrookly):

Evaluate tan2a if tan a=-5/12, pi/2< a< pi

OpenStudy (kkbrookly):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

see page 2 of http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf You will use this identity \[\Large \tan(2a) = \frac{2\tan(a)}{1-\tan^2(a)}\]

jimthompson5910 (jim_thompson5910):

let me know if that is enough to get started

OpenStudy (kkbrookly):

Would I start with 2tan(-5/12)/1-tan^2(-5/12)?

jimthompson5910 (jim_thompson5910):

it's not tan(-5/12) it's tan(a) = -5/12

jimthompson5910 (jim_thompson5910):

so everywhere you see a "tan(a)", replace it with -5/12

jimthompson5910 (jim_thompson5910):

\[\Large \tan(2a) = \frac{2\tan(a)}{1-\tan^2(a)}\] \[\Large \tan(2a) = \frac{2*\left(-\frac{5}{12}\right)}{1-\left(-\frac{5}{12}\right)^2}\] \[\Large \tan(2a) = ??\]

OpenStudy (kkbrookly):

-120/119?

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

I'm getting the same result

OpenStudy (kkbrookly):

Sweet. So that's the final answer?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (kkbrookly):

Thank you!!